OpenStudy (anonymous):

if H is 34 what is 1/2 h + 5 = ?

OpenStudy (accessdenied):

I am assuming H and h are the same and it is (1/2)*h... (1/2)(34) + 5 = 17 + 5 = 22

OpenStudy (anonymous):

22